A 40 foot beam is to be cut in two. If one portion is one third the length of the other portion, what are the two lengths of the portions?
Answer by edjones(8007)

Let x be the length of the other portion.
x/3 + x = 40
x+3x=120 multiply each side by 3.
4x=120
x=30'
x/3=10'
.
Ed